# Exploring Collaborative Immersive Visualization & Analytics for High-Dimensional Scientific Data through Domain Expert Perspectives

### v1

@@ Abstract.
This paper has been accepted to the Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26) and is shared here as a preprint version. The final, archival version will be available at https://doi.org/10.1145/3772318.3791203.
@@
copyright: nonejournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791203isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 4. Findings
4.1. Current Visualization & Analytics Ecosystem
Synthesized from interview findings, Figure 5 illustrates the workflow practices of domain experts in visualizing and analyzing high-dimensional scientific data, and Table 1 summarizes their primary analysis and visualization tools.
4.1.1. Data acquisition & storage infrastructure.
Domain experts report acquiring data from heterogeneous sources, ranging from structured databases and unstructured streams to satellite imagery, LiDAR scans, camera-trap photos, audio recordings, scientific waveforms, medical imaging systems, and real-time HPC telemetry. Several note infrastructures, where REDCap and PACS manage clinical and imaging data across numerous projects. Storage practices span local HDDs or SSDs for initial storage, cloud services like OneDrive and Google Drive for backup, and HPC clusters or AWS with Kubernetes for large-scale processing, often supported by distributed frameworks such as Dask. Experts require navigating frequent format transitions—moving from CSV for quick inspection to Parquet or Arrow for efficient processing, or to databases like DocDB for high-dimensional search, and one participant remarks that “Every format switch brings metadata inconsistencies and tool-specific constraints.” [P11]
4.1.2. Data preprocessing
4.1.3. Dimensionality reduction & embedding
4.1.4. Statistical analysis
4.1.5. Visualization design
appears in participants’ workflows as the step where high-dimensional analyses are translated into interpretable visual structures using scientific, statistical, and web-based tools. Several rely on ParaView, noting “…easier to load mesh data and apply filters like smoothing or point-flagging.” [P20] when working with 3D, LiDAR, or hierarchical structures. Python tools such as Matplotlib, Altair, and Vega-Lite support plots, heatmaps, and correlation matrices, while high-performance workflows pair with Polars for large DataFrame operations. Domain-specific representations include unit-sphere views for rotational motion, velocity maps, and 2D heatmaps, as noted, “Still, it’s really hard to show 3D visualizations” [P8]. Interactive displays often use D3.js and Observable for richer customization, and Stata, R, MATLAB, Tableau, and Plotly provide publication-ready figures and stakeholder dashboards for reporting.
4.1.6. Immersive & advanced visualization
4.1.7. Quality control
surfaces in participants’ practices as the process of ensuring computed results align with raw data and expected outcomes. Experts emphasized contextualizing outputs against raw inputs, noting the need to “show the computational results in the context of the raw data for QC.” [P16]. When multimodal data lack precise registration, participants rely on manual inspection, explaining that “most people just do side-by-side views and let the human eye take care of alignment” [P15]. QC routines include reproducibility checks by external and collaborative inspection through TensorBoard. Supporting infrastructure spans hierarchical data models, standardized web portals, and pseudonymized medical workflows, with domain-specific tasks using annotation tools to verify whether outputs match expected results.
4.1.8. Application-specific implementation
4.2. Current Collaboration Infrastructure
4.2.1. Team communication
relies on platforms such as Zoom or Microsoft Teams for synchronous discussion and Slack or email for asynchronous coordination, yet these remain disconnected from analytical environments. Participants struggled to convey insights about high-dimensional data through text-based channels, noting that “When sharing static screenshots, it’s very hard to explain my perspective - they can’t interact with the data” [P17]. This disconnect fragments workflows and causes critical context to be lost. Another participant explained, “Most platforms don’t let us talk over the same visualization in real time, so we end up describing things in words, which creates confusion.” [P12].Experts are forced to translate complex visual patterns into verbal descriptions, increasing misinterpretation and prolonging collaboration.
4.2.2. Sharing & distribution methods
predominantly involve exporting static images via email or cloud storage, removing the interactivity vital for scientific analysis. As one participant noted, “I just export a PNG or PDF, but then the interactive aspect is gone, so my collaborators can’t explore the data themselves.” [P19]. File-based sharing through Google Drive or Dropbox adds further friction due to limited integration with domain-specific visualization tools. The asynchronous nature exacerbates the problem, with another expert explaining, “We send Excel sheets or CSVs over email, and by the time someone plots them, it’s already outdated compared to my version.” [P11]. Such practices force repeated format conversions, contributing to inefficiencies and data integrity issues. Some teams use Jupyter or Google Colab to rerun code, but they still lack synchronous manipulation.
4.2.3. Version control
currently relies on manual file naming and direct communication, offering no systematic means to track changes or maintain provenance. As one participant explained, “If two of us modify the same dataset, it’s impossible to track which visualization is the latest without asking directly.” [P16]. While some experiment with GitHub, its file-size limits and the scalability constraints of Git LFS make it unsuitable for high-dimensional datasets, pushing teams toward ad-hoc workarounds such as appending version numbers, an approach one expert described as “version control becomes nightmare with multiple files (v1, v2, v3).” [P18]. Some teams use shared HPC servers to edit the same script and verify results by rerunning analyses across environments, “We share our code with them, and they run the code in their environment” [P11], highlighting the inconsistency of provenance tracking.
@@ 4.4.2. P2:
Immersive visualization as a catalyst for collaboration captures participants’ belief that shared immersive space can strengthen collective reasoning. Immersion was described as creating “It will feel like the same workspace, but with the immersive sense of being together.” [P1], enabling collaborators to “better brainstorm to develop the intuitions together” [P10] and build “a real deep understanding of the problem” [P6]. Participants emphasized immersive environments support communication across levels of expertise, allowing scientists to present insights “much more beneficial to the end users and stakeholders” [P11] who may not follow traditional plots. Multi-user interaction was seen as central, where “Many people can interact with the data, annotate the region of interest, [and] contrast and compare between different selections” [P5]. Several envisioned low-friction collaboration as “a Google doc for 3D data” [P13], where remote and co-located participants contribute within a shared spatial workspace in real time without moving files across platforms.
@@ 4.4.3. P3:
Workflow integration challenges and defense of traditional tasks reflect participants’ doubts about the productivity value of immersive environments while defending established workflows. Several raised skepticism about productivity gains, by cautioning “[IA] could be counterproductive because it’s so immersive and so cool that people will get distracted.” [P10] Precision-demanding tasks were described as poorly suited to XR, where interactions are “very hard to interact with the data if we can only use our fingers” [P1] and raw data processing remains “…very hard to process raw data actually” [P2]. Participants drew firm boundaries around established practices, emphasizing that “programming and scripting is easiest on my laptop” [P18] and that preprocessing should remain automated through familiar pipelines—“I would prefer just to write a Python script and get them all changed” [P17]. These concerns underscore that CIVA must align with, rather than replace, entrenched computational workflows to avoid disrupting productivity.
@@ 4.5.2. F2:
Real-time synchronized visualization keeps teams aligned by continuously updating shared scenes as the underlying data evolves. Participants stressed the need for high-frequency streaming, where “there is no break in updating the visualizations, as everything has to be streaming-based” [P16], ensuring collaborators remain synchronized with ongoing scientific measurements. Temporal controls further enhance collaborative reasoning by letting teams inspect evolving signals, “show the time variance of the waveform, not just statics in image” [P17] and revisit prior states through replay features that allow teams to “record all your steps while you do the data analysis, and replay the whole process.” [P5] for validation or comparison.
@@ 4.6.2. R2:
Data privacy restrictions limit use of CIVA because participants expect that legal and institutional rules will restrict how sensitive data enters immersive platforms. In medical domains, data cannot leave local infrastructure, since “we are not by law allowed to upload the data somewhere” [P16]. Participants also anticipate barriers when collaborating across organizations with incompatible policies, such as national laboratories, government agencies, or private companies, as noted, “data is useless if we have to share it through third-party apps.” [P15]
@@ 4.6.3. R3:
Interaction and comfort challenges reduce usability because participants expect imprecise gesture-based interaction, text-entry difficulty, and ergonomic strain to limit effective analysis. They predict discomfort, noting that they “can’t be productive with eyestrain and headache.” [P9], and foresee that immersive manipulation may feel “very hard to interact with if we can only use our fingers.” [P1]. Such usability burdens raise concerns about sustained analytical work in CIVA.
@@ 6. Limitations and Future Work
Our recruitment strategy relied on remote interviews with domain experts working with high-dimensional data, which may introduce sample skew toward experienced researchers. Though sessions were remote and participants could not interact with immersive systems directly, their expertise in data analysis, visualization, and immersive technologies provided a strong basis for articulating collaboration needs in CIVA. To mitigate potential priming, all participants viewed the same design-probe video and were asked open-ended questions. We acknowledge our positionality as researchers in IA and its influence on interpretation. Throughout the findings, we explicitly discuss how participant expertise, team structures, and methodological choices shaped the collaborative dynamics identified in our analysis. This reflexive framing informs both the interpretation of findings and the development of design implications.
As these platforms rely on advanced hardware and distributed infrastructure, future work should examine sustainable development models, including open-source ecosystems and cross-institutional partnerships to support adoption in under-resourced environments such as non-profit research centers and state-run medical facilities. Building on the design insights identified in this study, future work can develop CIVA systems for structured evaluations of collaboration quality, task performance, and accessibility. Comparative studies across 2D, immersive, and hybrid environments, supported by shared benchmark datasets and standardized evaluation frameworks, will enable systematic assessment of collaborative workflows. Incorporating perspectives beyond visualization and immersive analytics will further enrich understanding of CIVA’s role across diverse scientific domains.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791203isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 4. Findings
4.1. Current Visualization & Analytics Ecosystem
Synthesized from interview findings, Figure 5 illustrates the workflow practices of domain experts in visualizing and analyzing high-dimensional scientific data, and Table 1 summarizes their primary analysis and visualization tools.  [^3] Technical tools, frameworks, software, and methods mentioned in this section are listed in the supplementary document.
4.1.1. Data acquisition & storage infrastructure.
Domain experts report acquiring data from heterogeneous sources, ranging from structured databases and unstructured streams to satellite imagery, LiDAR scans, camera-trap photos, audio recordings, scientific waveforms, medical imaging systems, and real-time HPC telemetry. Several note infrastructures, where REDCap and PACS manage clinical and imaging data across numerous projects. Storage practices span local HDDs or SSDs for initial storage, cloud services like OneDrive and Google Drive for backup, and HPC clusters or AWS with Kubernetes for large-scale processing, often supported by distributed frameworks such as Dask. Experts require navigating frequent format transitions—moving from CSV for quick inspection to Parquet or Arrow for efficient processing, or to databases like DocDB for high-dimensional search, and one participant remarks that “Every format switch brings metadata inconsistencies and tool-specific constraints” [P11].
4.1.2. Data preprocessing
4.1.3. Dimensionality reduction & embedding
4.1.4. Statistical analysis
4.1.5. Visualization design
appears in participants’ workflows as the step where high-dimensional analyses are translated into interpretable visual structures using scientific, statistical, and web-based tools. Several rely on ParaView, noting “…easier to load mesh data and apply filters like smoothing or point-flagging” [P20], when working with 3D, LiDAR, or hierarchical structures. Python tools such as Matplotlib, Altair, and Vega-Lite support plots, heatmaps, and correlation matrices, while high-performance workflows pair with Polars for large DataFrame operations. Domain-specific representations include unit-sphere views for rotational motion, velocity maps, and 2D heatmaps, as noted, “Still, it’s really hard to show 3D visualizations” [P8]. Interactive displays often use D3.js and Observable for richer customization, and Stata, R, MATLAB, Tableau, and Plotly provide publication-ready figures and stakeholder dashboards for reporting.
4.1.6. Immersive & advanced visualization
4.1.7. Quality control
surfaces in participants’ practices as the process of ensuring computed results align with raw data and expected outcomes. Experts emphasized contextualizing outputs against raw inputs, noting the need to “show the computational results in the context of the raw data for QC” [P16]. When multimodal data lack precise registration, participants rely on manual inspection, explaining that “most people just do side-by-side views and let the human eye take care of alignment” [P15]. QC routines include reproducibility checks by external and collaborative inspection through TensorBoard. Supporting infrastructure spans hierarchical data models, standardized web portals, and pseudonymized medical workflows, with domain-specific tasks using annotation tools to verify whether outputs match expected results.
4.1.8. Application-specific implementation
4.2. Current Collaboration Infrastructure
4.2.1. Team communication
relies on platforms such as Zoom or Microsoft Teams for synchronous discussion and Slack or email for asynchronous coordination, yet these remain disconnected from analytical environments. Participants struggled to convey insights about high-dimensional data through text-based channels, noting that “When sharing static screenshots, it’s very hard to explain my perspective - they can’t interact with the data” [P17]. This disconnect fragments workflows and causes critical context to be lost. Another participant explained, “Most platforms don’t let us talk over the same visualization in real time, so we end up describing things in words, which creates confusion” [P12]. Experts are forced to translate complex visual patterns into verbal descriptions, increasing misinterpretation and prolonging collaboration.
4.2.2. Sharing & distribution methods
predominantly involve exporting static images via email or cloud storage, removing the interactivity vital for scientific analysis. As one participant noted, “I just export a PNG or PDF, but then the interactive aspect is gone, so my collaborators can’t explore the data themselves” [P19]. File-based sharing through Google Drive or Dropbox adds further friction due to limited integration with domain-specific visualization tools. The asynchronous nature exacerbates the problem, with another expert explaining, “We send Excel sheets or CSVs over email, and by the time someone plots them, it’s already outdated compared to my version” [P11]. Such practices force repeated format conversions, contributing to inefficiencies and data integrity issues. Some teams use Jupyter or Google Colab to rerun code, but they still lack synchronous manipulation.
4.2.3. Version control
currently relies on manual file naming and direct communication, offering no systematic means to track changes or maintain provenance. As one participant explained, “If two of us modify the same dataset, it’s impossible to track which visualization is the latest without asking directly” [P16]. While some experiment with GitHub, its file-size limits and the scalability constraints of Git LFS make it unsuitable for high-dimensional datasets, pushing teams toward ad-hoc workarounds such as appending version numbers, an approach one expert described as “version control becomes nightmare with multiple files (v1, v2, v3)” [P18]. Some teams use shared HPC servers to edit the same script and verify results by rerunning analyses across environments, “We share our code with them, and they run the code in their environment” [P11], highlighting the inconsistency of provenance tracking.
@@ 4.4.2. P2:
Immersive visualization as a catalyst for collaboration captures participants’ belief that shared immersive space can strengthen collective reasoning. Immersion was described as creating “It will feel like the same workspace, but with the immersive sense of being together” [P1], enabling collaborators to “better brainstorm to develop the intuitions together” [P10] and build “a real deep understanding of the problem” [P6]. Participants emphasized immersive environments support communication across levels of expertise, allowing scientists to present insights “much more beneficial to the end users and stakeholders” [P11] who may not follow traditional plots. Multi-user interaction was seen as central, where “Many people can interact with the data, annotate the region of interest, [and] contrast and compare between different selections” [P5]. Several envisioned low-friction collaboration as “a Google doc for 3D data” [P13], where remote and co-located participants contribute within a shared spatial workspace in real time without moving files across platforms.
@@ 4.4.3. P3:
Workflow integration challenges and defense of traditional tasks reflect participants’ doubts about the productivity value of immersive environments while defending established workflows. Several raised skepticism about productivity gains, by cautioning “[IA] could be counterproductive because it’s so immersive and so cool that people will get distracted” [P10]. Precision-demanding tasks were described as poorly suited to XR, where interactions are “very hard to interact with the data if we can only use our fingers” [P1] and raw data processing remains “…very hard to process raw data actually” [P2]. Participants drew firm boundaries around established practices, emphasizing that “programming and scripting is easiest on my laptop” [P18] and that preprocessing should remain automated through familiar pipelines—“I would prefer just to write a Python script and get them all changed” [P17]. These concerns underscore that CIVA must align with, rather than replace, entrenched computational workflows to avoid disrupting productivity.
@@ 4.5.2. F2:
Real-time synchronized visualization keeps teams aligned by continuously updating shared scenes as the underlying data evolves. Participants stressed the need for high-frequency streaming, where “there is no break in updating the visualizations, as everything has to be streaming-based” [P16], ensuring collaborators remain synchronized with ongoing scientific measurements. Temporal controls further enhance collaborative reasoning by letting teams inspect evolving signals, “show the time variance of the waveform, not just statics in image” [P17] and revisit prior states through replay features that allow teams to “record all your steps while you do the data analysis, and replay the whole process” [P5] for validation or comparison.
@@ 4.6.2. R2:
Data privacy restrictions limit use of CIVA because participants expect that legal and institutional rules will restrict how sensitive data enters immersive platforms. In medical domains, data cannot leave local infrastructure, since “we are not by law allowed to upload the data somewhere” [P16]. Participants also anticipate barriers when collaborating across organizations with incompatible policies, such as national laboratories, government agencies, or private companies, as noted, “data is useless if we have to share it through third-party apps” [P15].
@@ 4.6.3. R3:
Interaction and comfort challenges reduce usability because participants expect imprecise gesture-based interaction, text-entry difficulty, and ergonomic strain to limit effective analysis. They predict discomfort, noting that they “can’t be productive with eyestrain and headache” [P9], and foresee that immersive manipulation may feel “very hard to interact with if we can only use our fingers” [P1]. Such usability burdens raise concerns about sustained analytical work in CIVA.
@@ 6. Limitations and Future Work
Our recruitment strategy relied on remote interviews with domain experts working with high-dimensional data, which may introduce sample skew toward experienced researchers. Though sessions were remote and participants could not interact with immersive systems directly, their expertise in data analysis, visualization, and immersive technologies provided a strong basis for articulating collaboration needs in CIVA. To mitigate potential priming effects, all participants viewed the same design-probe video and were asked open-ended questions. We acknowledge our positionality as researchers in IA and its influence on interpretation. Throughout the findings, we explicitly discuss how participant expertise, team structures, and methodological choices shaped the collaborative dynamics identified in our analysis. This reflexive framing informs both the interpretation of findings and the development of design implications.
As CIVA platforms rely on advanced hardware and distributed infrastructure, future work should examine sustainable development models, including open-source ecosystems and cross-institutional partnerships to support adoption in under-resourced environments such as non-profit research centers and state-run medical facilities. Building on the design insights identified in this study, future work can develop CIVA systems for structured evaluations of collaboration quality, task performance, and accessibility. Comparative studies across 2D, immersive, and hybrid environments, supported by shared benchmark datasets and standardized evaluation frameworks, will enable systematic assessment of collaborative workflows. Incorporating perspectives beyond visualization and immersive analytics will further enrich understanding of CIVA’s role across diverse scientific domains.